OpenStudy (anonymous):

Choose the correct product of (8x - 4)(8x + 4). 64x^2 - 16 64x^2 + 16 64x^2 - 64x + 16 64x^2 + 64x + 16

OpenStudy (anonymous):

@bjan last one!

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

A

OpenStudy (anonymous):

It's alright if you have more questions. I'm avail till 10

OpenStudy (anonymous):

Are you sure its A? Lol

OpenStudy (anonymous):

And where are you located at? Its almost 1 am here in FL. Lol

OpenStudy (anonymous):

oh i'm in san ramon, ca

OpenStudy (anonymous):

and yes it's A for sure

OpenStudy (anonymous):

if you have more questions its fine

OpenStudy (anonymous):

Okay cool. Thanks so much for all of your help! Have a blessed evening!

OpenStudy (anonymous):

u too. thnks for the medals

OpenStudy (anonymous):

Thanks for the assistance! Lol

OpenStudy (anonymous):

lol bye

OpenStudy (triciaal):

difference of 2 squares (a + b) (a -b) = a^2 - b^2 in the case a = 8 x and b = 4 (8 x)^2 - 4^2 = 64 x^2 - 16 as stated A